# Optimal Stopping in Latent Diffusion Models

### v1

@@ Abstract
We identify and analyze a surprising phenomenon of Latent Diffusion Models (LDMs) where the final steps of the diffusion can degrade sample quality. In contrast to conventional arguments that justify early stopping for numerical stability, this phenomenon is intrinsic to the dimensionality reduction in LDMs. We provide a principled explanation by analyzing the interaction between latent dimension and stopping time. Under a Gaussian framework with linear autoencoders, we characterize the conditions under which early stopping is needed to minimize the distance between generated and target distributions. More precisely, we show that lower-dimensional representations benefit from earlier termination, whereas higher-dimensional latent spaces require later stopping time. We further establish that the latent dimension interplays with other hyperparameters of the problem such as constraints in the parameters of score matching. Experiments on synthetic and real datasets illustrate these properties, underlining that early stopping can improve generative quality. Together, our results offer a theoretical foundation for understanding how the latent dimension influences the sample quality, and highlight stopping time as a key hyperparameter in LDMs.
@@ 1 Introduction
A pivotal advancement in the evolution of diffusion models is the introduction of the Latent Diffusion Model (LDM, Rombach et al., 2022). Instead of performing the computationally intensive diffusion process in the high-dimensional pixel space, LDMs first compress the data into a lower-dimensional latent space using a pretrained autoencoder (AE, Kingma & Welling, 2013). The diffusion steps then occur within this more manageable latent representation, significantly reducing computational requirements and training time without a meaningful loss of quality. Once the generative process is complete, a decoder maps the resulting latent vector back into a full-resolution image.
Recent research has shown that classical diffusion models excel at learning the geometric structure of low-dimensional data (see, for example, Li & Yan, 2024; Azangulov et al., 2024). In addition, prior work shows that the final steps of the diffusion are essential to contract the probability mass onto the data manifold (George et al., 2025). However, the fidelity of generated samples critically depends on the numerical stability of the stochastic differential equation (SDE) solver used in the backward diffusion process. One well-documented challenge in this method is the onset of numerical instability as the timestep $t$ approaches 0 (Song et al., 2021). This phenomenon is primarily caused by a vanishing signal-to-noise ratio, which makes the corresponding score function difficult to accurately estimate and causes the SDE to become stiff. To avoid this, in practice both the training objective and the inference-time integration are restricted to the interval $0,T-\delta]$ (Vahdat et al., 2021) for some small, non-zero stopping time, $\delta>0$.
By contrast, this suggests a key benefit of LDMs, which to our knowledge has not been explored in the literature so far: by relying on the autoencoder to reduce the dimensionality, the LDM provides an alternative mean to learn low-dimensional manifolds without relying on the last few steps. This intuition motivates the following hypothesis:
In latent diffusion models, the last diffusion steps do not improve, or even degrade, sample quality.
We find empirical evidence of this hypothesis by comparing samples from a LDM with those of a standard diffusion model directly trained in the pixel space, both trained on the dataset CelebA. In the case of an LDM, degradation in the last sampling steps is evidenced by a rising FID score, as illustrated in Figure 1. In contrast, this phenomenon, which happens much earlier in the diffusion process than potential numerical instabilities close to $T$, is absent in standard diffusion models. Visual inspection of the associated images confirms that their quality does not improve in the last steps of the LDM, contrarily to standard diffusion (see Figure 2). Our main contribution in this work is to provide a theoretical justification of this observation. To this aim, we analyze the phenomenon using Gaussian data and a linear autoencoder. This choice is deliberate, as this simplified setting already exhibits phenomena similar to the larger-scale evidence, while being analytically tractable, allowing us to rigorously demonstrate the effect of early stopping and dimension reduction.
Refer to caption Refer to caption
Figure 1: (left) FID-30k score of latent diffusion model on CelebA-HQ, with latent shape $64\times 64\times 3$. (right) FID-30k score of standard diffusion model (trained in pixel space) on CelebA64 ($64\times 64\times 3$).
@@
Figure 2: Samples generated with a latent diffusion model (LDM) and a pixel-space diffusion. In the LDM, the before-last sample is nearly denoised and indistinguishable from the final one, whereas in the pixel-space model stronger noise remains at that timestep. See Appendix D for more examples.
@@
Our contributions are as follows:
  • We propose a theoretical framework to analyze the impact of the last steps of the LDM by studying the evolution, along the backward diffusion, of the Wasserstein-2 distance between the data distribution and the generated distribution for Gaussian data (Section 3).
  • We identify the optimal latent dimension as a function of the covariance matrix and the stopping time, for the case of a diagonal data covariance matrix. In particular, when the data lies on a linear subspace, we prove that the optimal strategy is to project the diffusion onto this subspace, and we determine the corresponding optimal stopping time. This reveals a time-dependent trade-off: distributions generated in the early stages of the backward process are best approximated in lower-dimensional spaces, while higher dimensions are required to faithfully reconstruct the data in the final sampling steps (Section 4).
@@
  • We extend our analysis to general covariance matrices, showing that the interaction between early stopping and latent dimensionality persists beyond the independent setting (Section 6).
@@ 2 Related work
Focusing on a theoretical analysis of this phenomenon, Achilli et al. (2025) investigate the optimal stopping time for diffusion models under the assumption that the data is concentrated on a low-dimensional manifold, a concept formalized by the Hidden Manifold Model (Goldt et al., 2020). Closer to our contribution is the work of Hurault et al. (2025). They also investigate the scenario where the true data distribution is Gaussian. Their analysis focuses on learning the score function using SGD, and allows them to determine an optimal stopping time. However, the study of these authors is limited to the diffusion model and did not consider the two-stage architecture of LDMs. Furthermore, the relationship between the data dimension and the derived optimal stopping time remained unexplored in their findings. In contrast, our work directly investigates the influence of the latent dimension on the optimal stopping time by incorporating an autoencoder into the diffusion model framework. We also demonstrate the need of early stopping without discretization of the backward diffusion process.
@@ 3 Notations and problem setup
Let $p_{0}$ be an unknown distribution in $\mathbb{R}^{D}$. With a slight abuse of notation, we use in the following the same notation for a distribution and its density function. The goal of diffusion models is to generate new observations following $p_{0}$, given an i.i.d. sample $(X_{1},\ldots,X_{n})$ drawn from $p_{0}$. The mechanism is as follows. Given a final diffusion time $T>0$, a latent dimension $d\leq D$, a matrix $P\in\mathbb{R}^{d\times D}$, and a scalar function $w:~[0,T]\to$$\mathbb{R}$, the latent forward variance-preserving (VP) SDE (Song et al., 2020) is defined by
@@
where $\overrightarrow{W_{t}}$ is a standard $D$-dimensional Brownian motion. The role of the matrix $P$ is to perform linear dimension reduction. Two special cases are of interest: first, if $d=D$ and $P$ is the identity matrix, we recover the standard formulation of diffusion models. Second, if $P$ projects on the first few principal components of the sample covariance matrix, this amounts to performing principal component analysis (PCA, Jolliffe, 2002). This projection is equivalent to linear autoencoders (Plaut, 2018), and there exists a pseudo-inverse $P^{+}\in\mathbb{R}^{D\times d}$ which allows us to map sample back to $\mathbb{R}^{D}$.
@@
$dP\overleftarrow{X}_{t}=(w_{T-t}^{2}P\overleftarrow{X}_{t}+2w_{T-t}^{2}s_{P}(P\overleftarrow{X}_{t},T-t))dt+\sqrt{2w_{T-t}^{2}}dP\overleftarrow{W_{t}},\quad P\overleftarrow{X}_{0}\sim P_{\#}p_{T},$ (2)

### v2

@@ Abstract
We identify and analyze a surprising phenomenon of Latent Diffusion Models (LDMs) where the final steps of the diffusion can degrade sample quality. In contrast to conventional arguments that justify early stopping for numerical stability, this phenomenon is intrinsic to the dimensionality reduction in LDMs. We provide a principled explanation by analyzing the interaction between latent dimension and stopping time. Under a Gaussian framework with linear autoencoders, we characterize the conditions under which early stopping is needed to minimize the distance between generated and target distributions. More precisely, we show that lower-dimensional representations benefit from earlier termination, whereas higher-dimensional latent spaces require later stopping time. We further establish that the latent dimension interplays with other hyperparameters of the problem such as constraints in the parameters of score matching. Crucially, this framework suggests that the reconstruction quality of the autoencoder alone can serve as a proxy to estimate the potential performance of the full LDM. Experiments on synthetic and real datasets illustrate these properties, underlining that early stopping can improve generative quality. Together, our results offer a theoretical foundation for understanding how the latent dimension influences the sample quality, and highlight stopping time as a key hyperparameter in LDMs.
Machine Learning, ICML
@@ 1 Introduction
A pivotal advancement in the evolution of diffusion models is the introduction of the Latent Diffusion Model (LDM, Rombach et al., 2022). Instead of performing the computationally intensive diffusion process in the high-dimensional pixel space, LDMs first compress the data into a lower-dimensional latent space using a pretrained autoencoder (AE, Kingma & Welling, 2013). The diffusion steps then occur within this more manageable latent representation, significantly reducing computational requirements and training time without a meaningful loss of quality. Once the generative process is complete, a decoder maps the resulting latent vector back into a full-resolution image. The standard training pipeline for LDMs operates as a two-stage process: an AE is trained first to compress the data, followed by the training of a diffusion model within the resulting latent space.
While recent research has focused on optimizing AE training to yield better latent representations for diffusion (e.g., Kouzelis et al., 2025; Zhou et al., 2025), selecting the appropriate latent dimension remains a critical challenge. The dimensionality must be fixed during the initial AE training phase, yet identifying the dimension that yields the optimal tradeoff between generation quality and compute cost typically requires the resource-intensive process of training a subsequent LDM for each candidate.
Furthermore, distinct behavioral differences exist between pixel-space and latent-space diffusion. As shown in Figure 1, latent-generated images appear to visibly stabilize in the final diffusion steps, in contrast to pixel-space diffusion, where the final steps are crucial for removing remaining noise. This challenges the common assumption that LDMs yield the best samples at the final timestep, suggesting instead that early stopping may actually improve image quality, as the decoder may introduce high-frequency artifacts in the final stages.
@@
Figure 1: Samples generated with a latent diffusion model (LDM) and a pixel-space diffusion. In the LDM, the before-last sample is nearly denoised and indistinguishable from the final one, whereas in the pixel-space model stronger noise remains at that timestep. See Appendix E for more examples.
@@
Our goal in this paper is to investigate the optimal stopping time and latent dimensionality for LDMs. More specifically, our contributions are as follows:
  • Working within a Gaussian framework, we explicitly derive the Wasserstein-2 distance, which is equivalent to the Fréchet distance in the Gaussian setting, between the true data distribution and the backward diffusion. In particular, we highlight that the diffusion process is identical to injecting noise into the latent representations prior to decoding (Section 3).
  • We determine the optimal latent dimension and stopping time for diagonal and general covariances. In particular, for data on a linear subspace, we prove that projecting onto that subspace is optimal. This reveals a trade-off: early generation steps favor lower dimensions, while later steps require higher dimensions for faithful reconstruction (Section 4).
@@
  • Our theoretical analysis suggests the existence of an optimal interval for each LDM, which can be identified solely by examining the corresponding noisy AEs. Building on this insight, we conduct experiments on ImageNet-256 to demonstrate that these results also have predictive power in practical settings, i.e., we are able to predict the optimal stopping time and optimal interval for different LDMs by the FID curves of the corresponding version of noisy AEs. We train a diverse set of AEs and their corresponding LDMs, ensuring a controlled comparison by maintaining a consistent total parameter count across models. We then show that the Fréchet inception distance (FID) of the decoded noisy latent representations closely tracks that of the backward latent diffusion process. Notably, the FID curve of different LDMs cross at the same time as their corresponding version of noisy AEs (Section 6).
@@ 2 Related work
Focusing on a theoretical analysis of this phenomenon, Achilli et al. (2025) investigate the optimal stopping time for diffusion models under the assumption that the data is concentrated on a low-dimensional manifold, a concept formalized by the Hidden Manifold Model (Goldt et al., 2020). Closer to our contribution is the work of Hurault et al. (2025). They also investigate the scenario where the true data distribution is Gaussian. Their analysis focuses on learning the score function using SGD, and allows them to determine an optimal stopping time. However, the study of these authors is limited to the diffusion model and did not consider the two-stage architecture of LDMs. Furthermore, the relationship between the data dimension and the derived optimal stopping time remained unexplored in their findings. In contrast, our work directly investigates the influence of the latent dimension on the optimal stopping time by incorporating an autoencoder into the diffusion model framework. We also demonstrate the need of early stopping without discretization of the backward diffusion process. Finally, recent findings suggest that full decoding (at $t=0$) does not always yield the lowest FID (e.g., Jayasumana et al., 2024). While this previous work attributed this behavior to be an artifact of the FID metric, in this work we challenge the common assumption that LDMs yield the best samples at the final timestep, suggesting instead that early stopping may actually improve image quality.
@@ 3 Notations and problem setup
Let $p_{0}$ be an unknown distribution in $\mathbb{R}^{D}$. With a slight abuse of notation, we use in the following the same notation for a distribution and its density function. The goal of diffusion models is to generate new observations following $p_{0}$, given an i.i.d. sample $(X_{1},\ldots,X_{n})$ drawn from $p_{0}$. The mechanism is as follows. Given a final diffusion time $T>0$, a latent dimension $d\leq D$, a semi-orthogonal projection matrix $P\in\mathbb{R}^{d\times D}$, and a scalar function $w:~[0,T]\to$$\mathbb{R}$, the latent forward variance-preserving (VP)-SDE (Song et al., 2020) is defined by
@@
where $\overrightarrow{W_{t}}$ is a standard $D$-dimensional Brownian motion. The role of the matrix $P$ is to perform linear dimension reduction. Two special cases are of interest: first, if $d=D$ and $P$ is the identity matrix, we recover the standard formulation of diffusion models. Second, if $P$ projects on the first few principal components of the sample covariance matrix, this amounts to performing principal component analysis (PCA, Jolliffe, 2002). This projection is equivalent to linear autoencoders (Plaut, 2018), and there exists a pseudo-inverse $P^{+}\in\mathbb{R}^{D\times d}$ which allows us to map sample back to $\mathbb{R}^{D}$. We leave to future work to extend the analysis of this phenomenon to other noise schedules such as EDMs (Karras et al., 2022).
@@
$dP\overleftarrow{X}_{t}$ $=(w_{T-t}^{2}P\overleftarrow{X}_{t}+2w_{T-t}^{2}s_{P}(P\overleftarrow{X}_{t},T-t))dt$
$\quad+\sqrt{2w_{T-t}^{2}}dP\overleftarrow{W_{t}},\quad P\overleftarrow{X}_{0}\sim P_{\#}p_{T},$ (2)